OpenStudy (anonymous):

If dy dx equals cosine squared of the quantity pi times y over 2 and y = 1 when x = 0, then find the value of x when y = 3.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

rearrange the equation: \[\frac{ dy }{ \cos ^{2} \left( \frac{ \pi y }{ 2 } \right) } = dx\] \[\int\limits_{}^{}\sec ^{2}\left( \frac{ \pi y }{ 2 } \right)=\int\limits_{}^{}dx\]

OpenStudy (anonymous):

Closed it by accident please ignore.

OpenStudy (anonymous):

This should look familiar to you

OpenStudy (anonymous):

Solve the differential equation using the integral I have given you and you will be able to compute the values of y and x and then plug in the values given in the problem.

OpenStudy (anonymous):

Just a question, what happened to dy?

OpenStudy (anonymous):

there should be a dy

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Not sure if I'm doing it right.

OpenStudy (anonymous):

I got \[\tan(y \pi/2) = x\]

OpenStudy (anonymous):

yes that is what I got too, but looking at it now, I see it doesn't satisfy the condition when y=1, x =0. Let me check and see where the error could be.

OpenStudy (anonymous):

Ok, so we're still on the same page.

OpenStudy (anonymous):

but also it is multiplied by the constant 2/pi which you get from u-substitution

jimthompson5910 (jim_thompson5910):

I'm getting \[\Large \frac{2}{\pi}\tan\left( \frac{\pi y}{2}\right) = x+C \] after integration. The problem is that when y = 1, the value of `tan(y*pi/2)=tan(1*pi/2)=tan(pi/2)` is undefined. So that's strange how they say `y = 1 when x = 0`

OpenStudy (anonymous):

@jim_thompson5910 that's the same thing I'm getting. Could be an error with the problem

OpenStudy (anonymous):

So how would I go about this... it's an automated system so I can't just tell em it doesn't make sense.

OpenStudy (anonymous):

Should I just play along?

jimthompson5910 (jim_thompson5910):

You'll have to bring it up with your teacher

jimthompson5910 (jim_thompson5910):

maybe they want you to use Euler's method here?

OpenStudy (anonymous):

Ypu could try to use that equation and plug in the value y=3 and see what you get, but I'm pretty sure the problem has an error at least with the initial condition y=1, x=0

OpenStudy (anonymous):

kind of going to be difficult though I think to beat the computer on this one, since you need to solve for C using the initial values, and you have an option of none of these, so you won't be able to be compltetely sure.

OpenStudy (anonymous):

I'll think I'm gonna go with none of these, thanks for pointing that out. When I plugged in 3 I got an extremely large number.

OpenStudy (anonymous):

@jim_thompson5910 what do you think?

jimthompson5910 (jim_thompson5910):

see attached for the slope field of dy/dx = cos^2(pi*y/2) If y = 1, then dy/dx = 0 Using euler's method, we'd be stuck on the red line of y = 1 no matter what x is. So it appears that y = 3 isn't possible to achieve. So I agree with @VeritasVosLiberabit I would say "none of these" and consult the teacher. This problem seems a bit off. Perhaps there's a typo somewhere? Then again, since y = 1 makes the f(x,y) function undefined, it is more valid to think of that red line as an asymptote if anything

OpenStudy (anonymous):

@VeritasVosLiberabit @jim_thompson5910 Thanks guys, much appreciated.